OpenStudy (anonymous):

Complete the following ordered pairs for the equation. 4x + 5y = -22 (_, -2) (2, _) (7, _)

OpenStudy (anonymous):

first one, replace \(y\) by \(-2\) and solve for \(x\) in the equation \[4x+5\times (-2)=-22\]

OpenStudy (anonymous):

you get \[4x-10=-22\] \[4x=-12\] \[x=-12\div 4=-3\]

OpenStudy (anonymous):

for the other two, replace \(x\) by the given number and solve for \(y\)

OpenStudy (anonymous):

okay thanks!!! :)